OpenStudy (anonymous):

Help please What is the percentage of nitrogen in N2O4? Show all calculations leading to an answer.

OpenStudy (anonymous):

To find the answer take the mass of all of the nitrogen atoms in the compound, divide by the molar mass, then multiply by 100 to express the answer as a percent. \[\color{green}{\frac{2 \times 14.0}{(2 \times 14.0) + (4 \times 16.0 )} \times 100}\]

OpenStudy (anonymous):

around 30% ?

OpenStudy (anonymous):

Closer to 29.2 % (3 significant figures)

OpenStudy (anonymous):

could you explain that? Just so I have a clear understanding. Thanks.

OpenStudy (anonymous):

I know about significant figures but in this equation im having a hard time figuring it out